Solve for the roots of the quadratic equation:
(3x^2-7x)/2=3
I would try to show work....but i don't know where to start...please help...i need to know this for a test tomorrow!!!!

Answer by scott8148(6628)

multiplying by 2 __ 3x^2-7x=6 __ subtracting 6 __ 3x^2-7x-6=0

factoring __ (3x+2)(x-3)=0

3x+2=0 __ subtracting 2 __ 3x=-2 __ dividing by 3 __ x=-2/3

x-3=0 __ adding 3 __ x=3